# Comparison results for Markov tree distributions

### v1

@@ 5.2 A special property of star structures
    The Schur order for conditional distributions has the interesting property that it implies an ordering for various well-known dependence measures, where a dependence measure is a functional $(X,Y)\mapsto\kappa(Y|X)$ such that (i) $\kappa(Y|X)\in[0,1]\,,$ (ii) $\kappa(Y|X)=0$ if and only if $X$ and $Y$ are independent, and (iii) $\kappa(Y|X)=1$ if and only if $Y$ is completely dependent on $X\,.$ For example, Chatterjee’s rank correlation, which has recently attracted a lot of attention in the statistical literature [10, 30, 41], is consistent with the Schur order for conditional distributions [4]. Due to Proposition 5.4, the Schur order also implies general comparison results with respect to the supermodular order. Roughly speaking, large elements in the Schur order for conditional distributions lead to strong dependencies and, if the dependencies are positive, to large elements in supermodular order.
Acknowledgement
The first author gratefully acknowledges the support of the Austrian Science Fund (FWF) project P 36155-N ReDim: Quantifying Dependence via Dimension Reduction and the support of the WISS 2025 project ’IDA-lab Salzburg’ (20204-WISS/225/197-2019 and 20102-F1901166-KZP).
Appendix
Appendix A Examples
The following example justifies the $\mathrm{\text{MTP}_{2}}$ assumption in Theorem 3.6 by showing that bivariate CI specifications do generally not lead to a Markov realization that is CI.
Example A.1 (A non-CI Markov tree distribution with bivariate CI marginals)
$P[X_{i}=k]$ $=1/3\quad\text{for }i,k\in\{0,1,2\}\,,$ (39)
$P[X_{i}=k,X_{i+1}=\ell]$ $=a^{i}_{k\ell}\quad\text{for }i\in\{0,1\}\text{ and }k,\ell\in\{0,1,2\}\,.$ (40)
Note that the univariate distributions in (39) are the marginals of the bivariate distributions in (40). From the definition of the matrices $a^{1}$ and $a^{2}$ it can be seen that the subvectors $(X_{0},X_{1})$ and $(X_{1},X_{2})$ are CI . It even follows that $(X_{0},X_{2})$ is CI , because, as a consequence of Lemma B.3, $X_{0}\uparrow_{st}(X_{1},X_{2})$ and $X_{2}\uparrow_{st}(X_{1},X_{0})\,.$ However, the vector $(X_{0},X_{1},X_{2})$ is not CI because
The next example shows that Theorem 3.7 fails to hold when the continuity assumption on the marginal distribution functions is not imposed. In addition, the example emphasizes that the dependence structure of a random vector determined by a Markov realization of a bivariate tree specification is not only determined by the set of bivariate copulas and the conditional independence structure. It also depends on the choice of the marginal distributions.
Example A.2 (General marginals are not sufficient for Theorem 3.7)
$F_{0}=F_{1}=F_{2}=G_{1}=G_{2}=F_{U(0,1)}\quad\text{and}\quad G_{0}=\mathds{1}_{[1,\infty)}.$
Then $F_{0},F_{1},F_{2},G_{1}$ and $G_{2}$ are continuous distribution functions and $G_{0}$ is the distribution function of the Dirac distribution in $1\,,$ which is not continuous. Denote by $M^{2}$ the bivariate upper Fréchet copula, i.e., $M^{2}(u,v)=\min\{u,v\}$ for $(u,v)\in[0,1]\,,$ and consider the bivariate specifications $B_{e}:=C_{e}:=M^{2}$ for $e\in E\,.$ Let $U$ and $V$ be independent and uniformly on $(0,1)$ distributed random variables. Then, for $F=(F_{0},F_{1},F_{2}),G=(G_{0},G_{1},G_{2})\,,$ $B=(B_{e})_{e\in E}\,,$ and $C=(C_{e})_{e\in E}\,,$ the random vectors $X=(X_{0},X_{1},X_{2}):=(U,U,U)$ and $Y=(Y_{0},Y_{1},Y_{2}):=(1,U,V)$ are Markov realizations of $(F,T,B)$ and $(G,T,C)\,,$ respectively, i.e.,
It holds that $F_{i}\leq_{lo}G_{i}$ for all $i\in\{0,1,2\}$ and $B_{e}=C_{e}$ (which trivially implies $B_{e}\leq_{lo}C_{e}$) for $e\in E\,.$ Further, the bivariate copulas $B_{e}$ and $C_{e}\,,$ $e\in E\,,$ are CI. Hence, all assumptions of Corollary 3.7i except of continuity of the marginal distributions are satisfied. However, it holds for all $(u,v)\in(0,1)$ that
$F_{X_{1},X_{2}}(u,v)=\min\{u,v\}>uv=F_{Y_{1},Y_{2}}(u,v)\,,$
which implies $(X_{1},X_{2})\nleq_{lo}(Y_{1},Y_{2})\,.$ Since the lower orthant order is closed with respect to marginalization, see, e.g., [87, Theorem 3.3.19], it follows that $X\nleq_{lo}Y$.
Replacing in the above setting $G_{0}$ by $G_{0}=\mathds{1}_{[0,\infty)}$ and considering $Y=(0,U,V)\,,$ it follows similarly that $(X_{1},X_{2})\nleq_{uo}(Y_{1},Y_{2})$ and thus $X\nleq_{uo}Y\,.$ Hence, we conclude that the continuity assumption for the marginal specifications in Corollary 3.7 cannot be omitted.
Note that in this example, $X_{1}$ and $X_{2}$ are comonotonic while $Y_{1}$ and $Y_{2}$ are independent. Since the bivariate tree specifications $(F,T,B)$ and $(G,T,C)$ only differ in the first component of the marginal specifications (i.e., $F_{0}\neq G_{0}\,,$ $F_{1}=G_{1}\,,$ $F_{2}=G_{2}\,,$ and $B=C$), we conclude that also the marginal distributions can affect the dependence structure of Markov tree distributions. In other words, conditional independence is not only a copula-dependent property.

### v3

@@ 4 A special property of star structures
    The Schur order for conditional distributions has the fundamental property that it implies an ordering of various dependence measures. A dependence measure is a functional $(X,Y)\mapsto\kappa(Y,X)$ such that (i) $\kappa(Y,X)\in[0,1],$ (ii) $\kappa(Y,X)=0$ if and only if $X$ and $Y$ are independent, and (iii) $\kappa(Y,X)=1$ if and only if $Y$ is perfectly dependent on $X.$ For instance, Chatterjee’s rank correlation, which has recently attracted a lot of attention in the statistics literature [11, 19, 27], is consistent with the Schur order for conditional distributions; see [3]. By Proposition 4.2, the Schur order also implies comparison results with respect to the supermodular order. Roughly speaking, large elements in the Schur order for conditional distributions lead to strong dependencies and, if the dependencies are positive, to large elements in supermodular order.
5 Distributional robustness in hidden Markov models
In this section, we apply our comparison results to hidden Markov models (HMMs), which form a subclass of Markov tree distributions. Using Theorem 1.3 and the results presented in Section 3 allows us to study distributional robustness of various functionals under ambiguity. As an illustrative example, we consider the maximum of a perturbed random walk and derive uncertainty bands for its distribution function in a setting with model uncertainty. For a comprehensive overview of hidden Markov models, including the subsequent definitions, we refer to [18, 30], along with the literature referenced therein.
Any HMM $(X,X^{*})$ has a functional representation, known as a (general) state-space model, by
for some measurable functions $f_{n},f^{*}_{n}\colon\mathbb{R}^{2}\to\mathbb{R}$ and i.i.d. uniformly on $[0,1]$ distributed random variables $\{\delta_{n}\}_{n\in\mathbb{N}},\{\varepsilon_{n}\}_{n\in\mathbb{N}_{0}},$ that are independent of the initial random variable $X_{0}$. In the language of Markov tree distributions, a HMM can equivalently be written as a sequence $Z=(Z_{n})_{n\in\mathbb{N}_{0}}$ of random variables defined by $Z_{2n}=X_{n}$ and $Z_{2n+1}=X^{*}_{n},$ for all $n\in\mathbb{N}_{0}$, where $Z$ follows a Markov tree distribution with respect to the tree
$T=(N,E),\text{ with }N=\mathbb{N}_{0}\text{ and }E=\big\{(i,j)\in 2\mathbb{N}_{0}\times\mathbb{N}_{0}\,|\,j-i\in\{1,2\}\big\};$ (13)
see Figure 3 for an illustration of the underlying tree structure.
There are two perspectives on the interpretation and application of HMMs. First, in fields such as communication theory, one can view the hidden process $X$ as a signal transmitted via a communications channel. Given the inherent noise in the channel, the receiver perceives the distortion $X^{*}$ and aims to reconstruct the original signal $X$; see e.g. [30, 44]. Second, in many models, such as in finance [52, 65], economics [36], climatology [1] or reliability theory [71], the focus lies directly on the observable process $X^{*}$, which is driven by a (latent) factor process $X$, the hidden process. For instance, $X^{*}$ may describe the market price of a stock, while $X$ models an economic factor influencing price fluctuations. In particular, any ARMA process has a representation of the form (12) with linear functions $f$ and $f^{*};$ see [2]. In the sequel, we adopt the latter perspective, focusing on drawing inferences from the observations of the hidden process.
As an interesting quantity, we consider the maximum of the observations $X_{0}^{*},\ldots,X_{n}^{*},$ noting that the maximum of a stochastic process is closely related to first passage times. Specifically, the maximum process exceeds a threshold $s$ at a certain time $n$ if and only if the first time where the underlying stochastic process enters the interval $(s,\infty)$ is less or equal than $n.$ For an overview of recent literature and for various applications related to first passage times and the maximum of a perturbed random walk, we refer [9, 10, 39, 40, 41, 42].
We make use of our comparison results to establish distributional robustness of $\max_{0\leq k\leq n}\{X_{k}^{*}\}$ in stochastic order under ambiguity of $X^{*}.$ Our findings are illustrated through uncertainty bands for the associated distribution function $F_{\max_{0\leq k\leq n}\{X_{k}^{*}\}}.$ Key is that the function
$f\colon(x_{0},\ldots,x_{n})\mapsto\mathds{1}_{\{\max\{0,x_{0},\ldots,x_{n}\}\leq t\}}$ (14)
is decreasing supermodular. As a direct consequence of Theorems 1.3, 3.1 and 3.5, the following result enables the comparison of various functionals (as outlined in Table 1) within classes of hidden Markov models.
Corollary 5.1 (Comparison results for hidden Markov models)
Let $(X,X^{*})$ and $(Y,Y^{*})$ be HMMs. Assume that $X_{n}^{*}\uparrow_{st}X_{n}$ for all $n\in\mathbb{N}$ and that $X_{n+1}\uparrow_{st}X_{n},$ $Y_{n}\uparrow_{st}Y_{n+1}$ as well as $Y_{n}^{*}\uparrow_{st}Y_{n}$ for all $n\in\mathbb{N}_{0}.$
  1. (i)
  2. (ii)
    Assume that $\overline{\mathsf{Ran}(F_{X_{n}})}=\overline{\mathsf{Ran}(F_{Y_{n}})}$ and $\overline{\mathsf{Ran}(F_{X_{n}^{*}})}=\overline{\mathsf{Ran}(F_{Y_{n}^{*}})}$ for all $n\in\mathbb{N}_{0}.$ If $C_{X_{n},X_{n+1}}\leq_{lo}C_{Y_{n},Y_{n+1}},$ $C_{X_{n},X_{n}^{*}}\leq_{lo}C_{Y_{n},Y_{n}^{*}},$ $X_{n}\leq_{st}Y_{n}$ (resp. $\geq_{st}$), and $X_{n}^{*}\leq_{st}Y_{n}^{*}$ (resp. $\geq_{st}$) for all $n\in\mathbb{N}_{0},$ then $(X,X^{*})\leq_{ism}(Y,Y^{*})$ (resp. $\leq_{dsm}$).
  3. (iii)
    Assume that all marginal distribution functions are continuous. If $C_{X_{n},X_{n+1}}\linebreak\leq_{lo}C_{Y_{n},Y_{n+1}}$ and $C_{X_{n},X_{n}^{*}}\leq_{lo}C_{Y_{n},Y_{n}^{*}}$ with $C_{Y_{n},Y_{n+1}}$ and $C_{Y_{n},Y_{n}^{*}}$ $\mathrm{\text{MTP}_{2}},$ and if $X_{n}\leq_{cx}Y_{n}$ and $X_{n}^{*}\leq_{cx}Y_{n}^{*}$ for all $n\in\mathbb{N}_{0},$ then $(X,X^{*})\leq_{dcx}(Y,Y^{*}).$
Remark 5.2
The SI assumptions in the above result are not very restrictive: Typically, there is a strong positive dependence between the latent variable $X_{n}$ and its perturbed observation $X_{n}^{*}.$ A standard example, where the SI condition $X_{n}^{*}\uparrow_{st}X_{n}$ is fulfilled, is the additive error model $f_{n}^{*}(x,z)=x+\sigma z$ for some $\sigma\geq 0.$ As with many models for stochastic processes, outcomes that are close in time exhibit strong positive dependence. Consequently, also the SI assumptions $X_{n+1}\uparrow_{st}X_{n}$ and $Y_{n}\uparrow_{st}Y_{n+1}$ are quite natural. Note that Corollary 5.1(i) can also be interpreted as an extension of Lemma 1.2 to a multi-factor model, where the hidden process serves as a common factor process.
[…]
Figure 3: The graphs illustrate the SI conditions on the hidden Markov processes $(X,X^{*})=(X_{n},X_{n}^{*})_{n\in\mathbb{N}_{0}}$ and $(Y,Y^{*})=(Y_{n},Y_{n}^{*})_{n\in\mathbb{N}_{0}}$ which lead to the comparison results in Corollary 5.1, where an arrow $U\to V$ indicates $V\uparrow_{st}U.$ The root of the underlying tree corresponds to the variable $X_{0}$ and $Y_{0},$ respectively. Note that only the SI condition between $X_{0}$ and $X_{0}^{*}$ can be dropped, see Proposition A.2.
5.1 A classical random walk
As a starting point, we model the hidden process as a classical random walk with independent and standard normal increments, i.e.,
$X_{n}=\sum_{i=1}^{n}\xi_{i},\quad X_{0}:=0,$ (15)
for i.i.d. $N(0,1)$-distributed random variables $\{\xi_{i}\}_{i\in\mathbb{N}}$. Then, $X=(X_{n})_{n\in\mathbb{N}_{0}}$ is a discrete-time Gauss-Markov process specified by the means $\mathbb{E}X_{n}=0$ and the covariances $\operatorname{Cov}(X_{i},X_{j})=\min\{i,j\}$ for all $i,j,n\in\mathbb{N}_{0}.$ Equivalently, by Proposition 2.4, $X$ can be considered as a discrete-time Markov process with marginal and bivariate copula specifications
$X_{n}$ $\sim N(0,n),$ (16)
$C_{X_{n},X_{n+1}}$ $=C_{\rho_{n,n+1}}^{\textnormal{Ga}},\quad\rho_{n,n+1}:=\mathrm{Cor}(X_{n},X_{n+1})=\sqrt{n/(n+1)},$ (17)
where $C_{\rho}^{\textnormal{Ga}}$ denotes the Gaussian copula with correlation parameter $\rho.$ For $\rho\geq 0,$ it is well known that $C_{\rho}^{\textnormal{Ga}}$ is CI, which is equivalent to
$X_{n+1}\uparrow_{st}X_{n}\quad\text{and}\quad X_{n}\uparrow_{st}X_{n+1}\quad\text{for all }n\in\mathbb{N}_{0}.$ (18)
The random walk $X$ is represented by the red paths in Figure 6, and the distribution function of the associated maximum $\max_{k\leq d}\{X_{k}\}$ is illustrated by the upper (red) graph in Figure 4 for $d=200.$
5.2 A perturbed random walk under model uncertainty
In the sequel, we model the observation process $X^{*}=(X_{n}^{*})_{n\in\mathbb{N}_{0}}$ by incorporating noise in various ways, and analyze how the noise affects the maximum observations.
Example 5.3 (Gaussian error model)
As a standard model for noise, we choose independent additive Gaussian errors to model the perturbed random walk $X^{*}=(X_{n}^{*})_{n\in\mathbb{N}},$ i.e.,
$X_{n}^{*}=X_{n}+\sigma_{n}\varepsilon_{n},\quad\sigma_{n}\in[0,\overline{\sigma}_{n}],$ (19)
for i.i.d. standard normal random variables $(\varepsilon_{i})_{i\in\mathbb{N}}$ independent of $X$ and for some deterministic $\overline{\sigma}_{n}>0.$ The parameter $\sigma_{n}$ is assumed to be ambiguous in the sense that it is only known to lie within the interval $0,\overline{\sigma}_{n}].$ For the family of models in (19), it can easily be verified that
$X_{n}^{*}$ $\sim N(0,n+\sigma_{n}),$ (20)
$C_{X_{n},X_{n}^{*}}$ $=C_{\rho_{n}}^{\textnormal{Ga}},\quad\text{with}\quad\rho_{n}:=\sqrt{n/(n+\sigma_{n})}.$ (21)
As a direct consequence of (19), the correlation of $(X_{n},X_{n}^{*})$ decreases with $\sigma_{n}$ and lies within the interval $\underline{\rho}_{n},1]$, for $\underline{\rho}_{n}:=\sqrt{n/(n+\overline{\sigma}_{n})}.$ We use our comparison results to show that, for fixed $d\in\mathbb{N},$ $\max_{n\leq d}\{X_{n}^{*}\}$ increases in $\sigma_{n}$ with respect to the stochastic order. To this end, let $Y_{n}^{*}:=X_{n}+\sigma_{n}^{\prime}\varepsilon_{n}$ for some fixed $\sigma_{n}^{\prime}\leq\sigma_{n}.$ Then we have
$F_{X_{n}^{*}}(x)$ $\geq F_{Y_{n}^{*}}(x)\quad\text{for all }x\leq 0,$ (22)
$F_{X_{n}^{*}}(x)$ $\leq F_{Y_{n}^{*}}(x)\quad\text{for all }x\geq 0,$ (23)
$C_{X_{n},X_{n}^{*}}$ $\leq_{lo}C_{Y_{n},Y_{n}^{*}}$ (24)
for all $n\in\mathbb{N}.$ For (22) and (23), we use the fact that the univariate normal distribution with zero mean is increasing in its variance parameter with respect to the convex order. Note that (23) is equivalent to $\max\{X_{n}^{*},0\}\geq_{st}\max\{Y_{n}^{*},0\},$ see also Remark 3.2(c). For (24), we use that the Gaussian copula family is increasing in its parameter with respect to the lower orthant order; see e.g. [4, Table 5]. Since the function $f$ in (14) is decreasing supermodular, it follows from Corollary 5.1(ii) that
$P(\max_{0\leq n\leq d}\{X_{n}^{*}\}\leq t)$ $=\mathbb{E}[\mathds{1}_{\max_{n\leq d}\{X_{n}^{*},0\}\leq t}]$
$\leq\mathbb{E}[\mathds{1}_{\max_{n\leq d}\{Y_{n}^{*},0\}\leq t}]=P(\max_{0\leq n\leq d}\{Y_{n}^{*}\}\leq t)$
for all $t\in\mathbb{R},$ using (23), (24), and (18) as well as $X_{k}^{*}\uparrow_{st}X_{k}$ and $Y_{k}^{*}\uparrow_{st}Y_{k}$ for all $k\in\mathbb{N}_{0}.$ Hence, defining $\underline{X}^{*}=(\underline{X}_{n}^{*})_{n\in\mathbb{N}_{0}}$ and $\overline{X}^{*}=(\overline{X}_{n}^{*})_{n\in\mathbb{N}_{0}}$ by
$\underline{X}_{n}^{*}:=X_{n}\quad\text{and}\quad\overline{X}_{n}^{*}:=X_{n}+\overline{\sigma}_{n}\varepsilon_{n},$ (25)
we obtain $P(\max_{n}\{\underline{X}_{n}^{*}\}>t)\leq P(\max_{n}\{X_{n}^{*}\}>t)\leq P(\max_{n}\{\overline{X}_{n}^{*}>t\})$ for all $t.$ This implies distributional robustness of the extreme order statistics in stochastic order. While $\underline{X}^{*}$ represents the setting without error, the process $\overline{X}^{*}$ models the scenario with the largest observation errors, as determined by the parameter bounds $\overline{\sigma}_{n}.$ In Figure 4, we illustrate the uncertainty bands for $\overline{\sigma}_{n}=3$ (left plot) and $\overline{\sigma}_{n}=0.3n$ (right plot).
Figure 4: Uncertainty bands of the distribution function $t\mapsto F_{\max\{X_{0}^{*},\ldots,X_{d}^{*}\}}(t)$ in the classes of hidden Markov models considered in Theorem 5.4 and Examples 5.6 and 5.7, where $d=200$ and $\sigma_{n}=3$ (left plot) resp. $\sigma_{n}=0.3n$ (right plot). The upper (red) graph is the distribution function of $\max_{n\leq d}\{\underline{X}_{n}^{*}\}$, which corresponds to the hidden Markov model without observation error. The black/orange/blue graph corresponds to the distribution function of $\max_{n\leq d}\{\overline{X}^{*}_{n}\}$ in the Gaussian/Clayton/survival Clayton setting, considering maximal uncertainty in Example 5.3/5.6/5.7.
In Example 5.3, the hidden Markov model $(X,X^{*})$ follows a multivariate normal model, for which comparison results are well known; see [57]. However, the bounds $\underline{X}^{*}=(\underline{X}_{n}^{*})_{n\in\mathbb{N}}$ and $\overline{X}^{*}=(\overline{X}_{n}^{*})_{n\in\mathbb{N}}$ in (25) are applicable to a much more general setting involving the marginal and copula ambiguity sets
$\mathcal{F}_{n}$ $:=\{F\text{ continuous cdf}\mid F_{N(0,n)}(t)\geq F(t)\geq F_{N(0,n+\overline{\sigma}_{n})}(t)\penalty 10000\ \forall t\geq 0\},$ (26)
$\mathcal{C}_{n}$ $:=\{C\text{ bivariate copula}\mid C\text{ SI and }C\geq_{lo}C_{\underline{\rho}_{n}}^{Ga},\penalty 10000\ \underline{\rho}_{n}=\sqrt{n/(n+\overline{\sigma}_{n})}\}.$ (27)
Here, $F_{N(0,n)}$ is the distribution function of $\underline{X}_{n}^{*}$, $F_{N(0,n+\overline{\sigma}_{n})}$ is the distribution function of $\overline{X}_{n}^{*},$ and $C_{\underline{\rho}_{n}}^{Ga}$ is the copula of $(X_{n},\overline{X}_{n}^{*}).$ The following result summarizes these findings.
Theorem 5.4 (Distorted random walk under ambiguity)
Consider the processes $\underline{X}^{*}$ and $\overline{X}^{*}$ in (25). Then, for any hidden Markov model $(X,X^{*})=(X_{n},X_{n}^{*})_{n\in\mathbb{N}_{0}}$ specified by the random walk $X$ in (15) and by $F_{X_{n}^{*}}\in\mathcal{F}_{n}$ and $C_{X_{n},X_{n}^{*}}\in\mathcal{C}_{n}$ for all $n\in\mathbb{N},$ we have
$\max_{0\leq n\leq d}\{\underline{X}_{n}^{*}\}\leq_{st}\max_{0\leq n\leq d}\{X_{n}^{*}\}\leq_{st}\max_{0\leq n\leq d}\{\overline{X}_{n}^{*}\}.$ (28)
Remark 5.5
As an extension of the setting in (19), Corollary 5.1 is also applicable when the observation process is modeled as $X_{n}^{*}=f_{n}(X_{n},\varepsilon_{n})$ for some componentwise increasing function $f\colon\mathbb{R}^{2}\to\mathbb{R}.$ Then, $X_{n}^{*}$ is still SI in $X_{n}\,;$ we refer to Examples 5.6 and 5.7 for this modified setting. Moreover, it is also possible to model the hidden process under ambiguity. For instance, the parameter $\rho_{n,n+1}$ in (17) may be assumed to belong to a subinterval of $[0,1].$ Recall that the lower orthant order, the SI property, and copulas are invariant under increasing transformations; see [69, Theorem 6.G.3] and [17, Proposition 2.12 and Theorem 3.3]. Consequently, the setting in Theorem 5.4 also applies to transformed processes such as $(\exp(X_{n}^{*}))_{n\in\mathbb{N}_{0}}.$
In Example 5.3 above, the perturbations are modeled as additive errors independent of $X_{n}.$ In the sequel, we investigate the behavior of the observation process when the errors depend on the values of the hidden process $X_{n}.$ As an illustrative example, we use Clayton and survival Clayton copulas to specify the dependencies of $(X_{n},X_{n}^{*})$, where
$C_{\theta}^{\textnormal{Cl}}(u,v)$ $:=\max\{(u^{-\theta}+v^{-\theta}-1)^{-1/\theta},0\}\quad\text{ and }$ (29)
$\quad C_{\theta}^{\textnormal{SCl}}(u,v)$ $:=1-u-v+C_{\theta}^{\textnormal{Cl}}(u,v),\quad(u,v)\in[0,1]^{2}$ (30)
denote the Clayton and survival Clayton copula with parameter $\theta\in(0,\infty).$ For $\theta=0$ the Clayton and survival Clayton copula are defined as $C_{\theta}^{\textnormal{Cl}}(u,v)=C_{\theta}^{\textnormal{Cl}}(u,v)=uv$, representing independence. For $\theta\to\infty,$ they model comonotonicity. While Clayton copulas exhibit lower tail dependencies, their associated survival copulas exhibit upper tail dependencies; see e.g. [4] and Figure 5.
Figure 5: Samples of a Gaussian copula (left), Clayton copula (mid), and survival Clayton copula (right), each having Kendall’s tau value $\tau=0.795$ which corresponds to the parameter $\rho=\sqrt{9/10}$ for the Gaussian copula and $\theta=7.764$ for the Clayton and survival Clayton copula. The plots indicate that the Clayton copulas exhibit lower tail-dependencies, while the survival Clayton copulas have upper tail-dependencies.
Example 5.6 (Clayton copula error model)
To analyze the influence of lower tail dependencies on $\max_{k\leq n}\{X_{k}^{*}\}$, we now consider Clayton copulas as dependence specifications of the observations process. The hidden process $X$ is modeled as before by the random walk in (15). For better comparability with the setting in Example 5.3, we assume that $X^{*}$ is specified by the same marginals as in the Gaussian setting (see (20)). However, $X_{n}^{*}$ and $X_{n}$ are now coupled by a Clayton copula, i.e.,
$X_{n}^{*}$ $\sim N(0,n+\sigma_{n}),\quad\sigma_{n}\in[0,\overline{\sigma}_{n}],$ (31)
$C_{X_{n},X_{n}^{*}}$ $=C_{\theta_{n}^{*}}^{\textnormal{Cl}},\quad\text{with}\quad\theta^{*}\in[\underline{\theta}_{n},\infty).$ (32)
For modeling a similar strength of positive dependence, we choose the lower bound for the parameter interval as $\underline{\theta}_{n}:=\theta(\underline{\rho}_{n})$, where
$\theta(\rho):=\frac{2\arcsin(\rho)}{\pi-2\arcsin(\rho)}$ (33)
is the parameter $\theta$ such that the Clayton copula $C_{\theta}^{\textnormal{Cl}}$ has the same Kendall’s tau value as the Gaussian copula with parameter $\rho\,;$ see e.g. [4, Table 6]. Recall that $\underline{\rho}_{n}$ is specified in Example 5.3. It is well known that any Clayton copula $C_{\theta}^{\textnormal{Cl}}$ is CI for every $\theta\geq 0$, and that the Clayton copula family is $\leq_{lo}$-increasing in its parameter; see e.g. [4, Table 3]. Thus, specifying $\underline{X}^{*}=(\underline{X}_{n}^{*})_{n\in\mathbb{N}}$ and $\overline{X}^{*}=(\overline{X}_{n}^{*})_{n\in\mathbb{N}}$ by
$\underline{X}_{n}^{*}$ $=X_{n},$ (34)
$F_{\overline{X}_{n}^{*}}$ $=F_{N(0,n+\overline{\sigma}_{n})}\quad\text{and}\quad C_{X_{n},\overline{X}_{n}^{*}}=C_{\underline{\theta}_{n}}^{\textnormal{Cl}},$ (35)
we can derive similar comparison results as in Theorem 5.4, where now the Gaussian copula in (27) is replaced by the Clayton copula $C_{\underline{\theta}_{n}}^{\textnormal{Cl}}.$ The uncertainty bands for the distribution function of $\max_{k\leq n}\{X_{k}^{*}\}$ are visualized in Figure 4.
Figure 6: Samples from the hidden process $X=(X_{n})_{n\in\mathbb{N}_{0}}$ and the observations process $\underline{X}^{*}=(\underline{X}^{*}_{n})_{n\in\mathbb{N}_{0}}$ for $n\in[0,200]$ in various settings: The red graphs represent two sample paths of the hidden process in (15) (which coincides with the observation process $\overline{X}^{*}$ in the absence of errors). The black/orange/blue graphs model the observation $\underline{X}^{*}$ in the Gaussian/Clayton/survival Clayton setting under maximal uncertainty, as specified in Example 5.3/5.6/5.7. For this plot, we set $\overline{\sigma}_{n}=3$ for all $n.$
As illustrated in Figure 6, $X_{n}^{*}$ exhibits a strong dependence on $X_{n}$ for small values of $X_{n}.$ Conversely, when $X_{n}$ attains large values, $X_{n}^{*}$ becomes less dependent on $X_{n}$, resulting in significant fluctuations. Consequently, in the Clayton copula setting, the uncertainty interval for $F_{\max_{k\leq n}\{X_{k}^{*}\}}(t)$ is wider for $t\gg 0$ (see Figure 4), leading to greater uncertainty in the first-passage times.
In the following example, we analyze the influence of upper tail dependencies between the hidden process and the observations. For simplicity, these dependencies are modeled using survival Clayton copulas.
Example 5.7 (Survival Clayton copula error model)
To analyze the influence of upper tail dependencies, we consider a setting similar to Example 5.6, replacing the Clayton copula with the survival Clayton copula defined in (30). By straightforward symmetry arguments, survival Clayton copulas are also CI and $\leq_{lo}$-increasing in their parameter. Hence, the setting from Example 5.6 similarly applies to survival Clayton copulas, allowing us to derive uncertainty bands for $F_{\max_{k\leq n}\{X_{k}^{*}\}}$ in an analogous way.
Figure 6 also illustrates the behavior of the perturbed random walk in the survival Clayton copula setting. Large values of $X_{n}$ are strongly coupled with $X_{n}^{*},$ whereas the observation process exhibits greater fluctuations for small values of $X_{n}.$ Consequently, there is less uncertainty in the distribution of $\max_{k\leq n}\{X_{k}^{*}\}$ (see Figure 4) and, therefore, also less ambiguity in the first-passage time process.
6 Counterexamples
In this section, we provide various examples that demonstrate the generality of our results.